What is the angular speed (in rad/s) of a body whose frequently is 50Hz?ANS:100pi rad/s.


1
Expert's answer
2021-03-01T12:40:02-0500
"\omega=2\pi f=2\pi\cdot50\ Hz=100\pi\ \dfrac{rad}{s}."